OpenStudy (anonymous):

\[ (\sqrt{2}+1 )^n=\sqrt{m-1}+\sqrt{m}\]

OpenStudy (anonymous):

\[\color{red}{\text{show that for constants m,n} (\sqrt{2}+1 )^n=\sqrt{m-1}+\sqrt{m}}\]

OpenStudy (experimentx):

is m,n natural or real number?

OpenStudy (anonymous):

Let m and n be are positive integer

OpenStudy (experimentx):

it seems to be valid for any number http://www.wolframalpha.com/input/?i=solve+%281+%2B+sqrt%282%29%29%5Epi+%3D+sqrt%28m+-+1%29+%2B+sqrt%28m%29

OpenStudy (anonymous):

yes i was supprised when i teseted 1,2,3... but the question says for some positive integer m.

OpenStudy (anonymous):

oh sorry its not pi its n

OpenStudy (experimentx):

interestingly i found the n=1,2,3, ... the value of m is given by http://oeis.org/search?q=2%2C9%2C50%2C+289%2C+1682&sort=&language=english&go=Search

OpenStudy (anonymous):

oh exaclty this is true since any triangular number is \[T=\frac{n^2-n}{2}\implies 2T=n^2-n \implies (n-1/2)^2=2T+1/4 \implies 2(n-1/2)=\sqrt{8T+1}\] stuck!

OpenStudy (anonymous):

but yeah thats true

OpenStudy (anonymous):

the equations not very clear they cud use a LaTex

OpenStudy (experimentx):

All Boils down to proving that `Table[Sum[Binomial[2 n, 2 k - 1] 2^{k - 2}, {k, 1, n}] , {n, 1, 10}]` is equal to `Table[Sqrt[m (m - 1)/2] , {m, {1, 2, 9, 50, 289, 1682, 9801, 57122}}]`

OpenStudy (experimentx):

\[ (\sqrt{2}+1)^n = a+ b \sqrt 2 \\ (\sqrt{2}-1)^n = -a+ b \sqrt 2 \\ 1 = 2 b^2 -a^2 \\ 2b^2 = a^2 - 1 \\ \] Follows that \[ (\sqrt{2} + 1)^n = \sqrt{a^2} + \sqrt{a^2 + 1}\]

OpenStudy (experimentx):

\[ \sum_{k=1}^n \binom{2n}{2k-1} 2^{k-2} = \sqrt{\frac{m(m+1)}{2}}\] is a consequence of the question. I doubt if it could be proved other way.